OpenStudy (anonymous):

how to solve dy/dt=1/2(sinx+cosx) differential equation?

OpenStudy (anonymous):

aany one please reply me?

OpenStudy (anonymous):

dy/dt with respect to x??

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

u can' solve it. Not when it's dy/dt and with respect to x

OpenStudy (anonymous):

can you please help me how to solve with respective to dt.?

OpenStudy (anonymous):

Integral please help me...

OpenStudy (anonymous):

kramen the question should be: dy/dt=1/2(sint+cost)

OpenStudy (anonymous):

okay. then which method is easy to solve the problem?

OpenStudy (anonymous):

Integral help me..

OpenStudy (anonymous):

Just integrate the whole thing that's all use: wolframalpha.com

OpenStudy (anonymous):

its a separable equation: \[\frac{dy}{dt}= \frac{1}{2}(sint+cost)\] \[dy=\frac{1}{2}(sint+cost)dt\] integrate both sides: \[\int\limits dy= \frac{1}{2} \int\limits (sint +cost)dt\] \[\int\limits dy= \frac{1}{2}[\int\limits sint dt+\int\limits costdt]\] \[y=\frac{1}{2}(sint-cost)+C\]

OpenStudy (anonymous):

Integral help me..

OpenStudy (anonymous):

thank you nadeem and integral..! have a gr8 day..